9. Write the solution to 2x²y +22xyt by=0 for L= 42, 50, 58?
Answer #1

m-2 Given 2x 7 +22xy +YL = o. L342,50,58 ? het dessume x - xm Y = mam; y = m(M-12 Substitue of, yy in Abore equation i 2x²-52 for L=50 2m²+ 20m +50=0 Roots m, = m₂ = -5 for dildo Game Roots m. = ma solution is (arxea) me = [, +25] e final station